# Enhancing Repository-Level Code Generation with Integrated Contextual Information

### v1

@@ 3.2.1. Type Context
Figure 3. The dependency graph of the Java class CMAESOptimizer
Figure 3 shows an undirected dependency graph $G=(V,E)$, in which each vertex corresponds to a type and each edge represents the dependency relationship. In particular, if type B occurs in the implementation of type A, there should be an edge between A and B, and vice versa. Thus, the dependency graph contains types that we are likely to use and refer to when writing code of method body.
In Figure 3, we build the dependency graph starting from class CMAESOptimizer. As the graph spans, the vertex set and the edge set will eventually converge, and we can obtain the types that are relevant to CMAESOptimizer. To pass such type information to the LLM, we create type context from the graph by extracting the fields and function signatures within the types and concatenating them into a string.
However, in large code repositories, the number of vertices could grow exponentially, making the resulting type context too long to fit in an LLM’s context window. In addition, higher-order neighbor vertices in the graph tend to have less relevance to each other. For the example in Figure 3, programmers are more likely to use the API of RealMatrix rather than RealVector, which is a higher-order neighbor of CMAESOptimizer compared to RealMatrix. Therefore, in practice, we limit the scope of type context creation to one-order neighbors in the graph. Assuming that LLMs have knowledge of the standard libraries of many languages, we also prune the types in these libraries.
@@ 3.2.2. Building Type Context via Static Analyzer
To get detailed information (i.e., fields and methods) of a type from its name, some straightforward heuristics might be used, e.g., locating imports or looking for similar file names. However, these heuristics are not sound and do not work with all statically typed programming languages. Thus, we need to perform an analysis to accurately navigate to a type and find its definitions and implementations. Many language servers serve as static analyzers and provide this type-based analysis, such as clangd (cla, 2024) for C++, Eclipse JDT (ecl, 2024) for Java, and rust-analyzer (rus, 2024b) for Rust.
Specifically, to build the type context for a function code generation task, we first determine the set of relevant types (as Section 3.2.1 describes). We start with the signature and form a set of types, and the initial element is the type to which the current function belongs. Then we implement a parser that works collaboratively with the static analyzer. The static analyzer provides detailed type information to the parser. The parser then searches for a new occurring type, adds it to the set, and queries the analyzer with its location (i.e., byte offset or line and column offsets in a file). This interaction ends when all one-order neighbors of the initial types are found. Finally, we concatenate the fields and method signatures within the set of types and obtain the type context.
@@ 3.3. Code Generation via LLM
Finally, we describe how to utilize the previous results and invoke the LLM to generate code. Figure 4 illustrates our prompt design, in which there are two fields: CONTEXT and CODE. To construct a prompt, we fill the CONTEXT field with type context and retrieved code sequentially, and place the docstrings and the signature of the method at the beginning of the CODE field. The prompt serves as input to a frozen LLM [^7] We discuss the details of selected models and model access in Section 4.4., and we expect the model to complete the CODE field by generating the method body (i.e., the green part in Figure 4). In addition, to post-process the raw output of the LLM, we employ several procedures [^8] The post-processing procedures may vary with different LLMs and programming languages., including truncating texts, prepending function signatures, and removing Markdown syntaxes.
@@ 4. Experimental Setup
In this section, we first formulate three research questions (RQs). Based on the RQs, we describe the details of the baselines, evaluation metric, benchmarks, and implementation of our approach. We aim to answer the following research questions:
  1. RQ.1
    How effective is our approach at generating function code in a repository, compared with the baselines?
  2. RQ.2
    How do the two components in our approach (i.e., relevant code retrieval and type context extraction) contribute to the overall effectiveness respectively?
  3. RQ.3
@@
In brief, we evaluate the effectiveness of our approach in RQ1, and then conduct an ablation study in RQ2. Finally, in RQ3 we study the generalizability of our approach and the performance across several LLMs. For implementation and evaluation, we select two statically typed programming languages, Java and Rust, since Java is one of the most widely used programming languages (tio, 2024), and Rust is a new system-level language that attracts the attention of many developers (rus, 2024c).
@@ 4.1. Baselines
As described in Section 3.3, Figure 4 shows the prompt template used in our evaluation. As the CONTEXT field is the only variable, we compare our approach with methods that fill different contexts in the prompt. For the repository-level code generation task, there are three types of context: (1) No context, (2) Local context, and (3) Cross-file context. Therefore, we design the following baselines for comparison in RQ1:
@@
    Vanilla: In this method, we prompt the LLMs without any repository context. In this case, all natural language descriptions are given, but the CONTEXT field in Figure 4 is left empty.
@@
    In-File: In this method, we consider code generation as a left-to-right file completion task. We extract code before the function to be generated, and put it in the CONTEXT field.
@@
    RepoCoder: RepoCoder (Zhang et al., 2023) is an approach that augments the prompt by iteratively retrieving referential code fragments, and it is originally proposed for Python code repositories. We adopt their idea and implement this baseline for Java and Rust based on our retrieval pipeline. Specifically, the initial iteration uses natural language descriptions (i.e., docstrings) as the query for code retrieval, and the second iteration uses previously generated code for retrieval.
4.2. Evaluation Metric
For function code generation task, we argue metrics based on text similarity are not accurate, and it is essential to use metrics based on execution results. Therefore, to evaluate the functional correctness of the generated code, we follow Chen et al. (Chen et al., 2021) and use the unbiased pass@$k$ as a metric. This metric is defined as
[…]
where $n$ is the number of code samples generated per task, $c$ is the number of functional correct code samples, and $k$ is the parameter in “pass@$k$” ($n\geq k$). Here, a code sample is functionally correct if and only if it passes the tests in the benchmark or dataset. We calculate pass@$k$ values for each task, and the final pass@$k$ for a benchmark is the mean value of all pass@$k$ values.
@@ 4.3.1. Java Benchmark
There are several publicly available repository-level benchmarks for Java (Ding et al., 2024; Yu et al., 2024; Agrawal et al., 2024; Zeng et al., 2024). However, it is challenging to use many of them to evaluate CatCoder because they either lack the raw repository code or are not designed for function code generation tasks. Considering the challenges, the CoderUJB (Zeng et al., 2024) benchmark is a suitable base benchmark that contains 238 code generation tasks, and we successfully adapt it for our evaluation. In the adaptation process, we filter out a few functions that satisfy the following patterns:
@@
    The length of function code exceeds max output limit (i.e., 512 tokens in our settings).
@@
    The function is standalone, which means that it can be moved to a new class [^9] We assume that all Java standard libraries are imported in advance when applying this rule. and still compiles. The removal of standalone functions helps better evaluate the performance of repository-level code generation.
@@
    The function does not have any modifiers and annotations. This helps us perform post-processing more easily.
After filtering, we finally obtain 199 coding tasks. As the methods in the benchmark are mined by the CoderUJB authors from the Defects4J (Just et al., 2014) dataset, it is convenient to test the correctness of the generated code utilizing the existing unit test methods. Specifically, we “patch” the repository with generated code and run the defects4j test -t ... command.
@@ 4.3.2. Rust Benchmark
To our knowledge, the only code generation benchmark for the Rust programming language is MultiPL-E (Cassano et al., 2023), which translates the original HumanEval (Chen et al., 2021) and MBPP (Austin et al., 2021) benchmarks. However, this benchmark is designed for standalone function code generation tasks and thus is unsuitable for our evaluation. In this case, we construct a new benchmark named RustEval.
@@
As shown in Table 1, we collect 13 repositories of various categories from crates.io (cra, 2024), the official Rust package registry. In these repositories, we obtain 90 functions with docstrings that satisfy the official guidelines (rus, 2024a) (i.e., have detailed descriptions and at least one documentation test). To test the correctness of the generated code, we utilize the documentation tests in the docstrings by running the cargo test --doc ... command.
@@ 4.4.1. Approach
We implement CatCoder for Java and Rust, and the source code is written in Python and Rust. Specifically in Relevant Code Retrieval, we use MPNet V2 (mpn, 2024) as the embedding model. It is an open-source model fine-tuned from the base MPNet (Song et al., 2020) model using massive training data, including the popular CodeSearchNet (Husain et al., 2019). In Type Context Extraction, we use Eclipse JDT.LS (ecl, 2024) and rust-analyzer (rus, 2024b) as the static analyzers. For Java, our approach interacts with Eclipse JDT.LS through the Language Server Protocol (lsp, 2024). For Rust, we use rust-analyzer as a third-party library and implement a binding between Rust and Python.
@@ 4.4.2. Model Selection
(Roziere et al., 2023) CodeLlama-7b-Instruct CL-7B
(Roziere et al., 2023) CodeLlama-13b-Instruct CL-13B
(Roziere et al., 2023) CodeLlama-34b-Instruct CL-34B
(Guo et al., 2024) deepseek-coder-6.7b-instruct DS-6.7B
(cod, 2024) codegemma-7b-it CG-7B
(Bai et al., 2023) CodeQwen1.5-7B-Chat CQ-7B
(Lozhkov et al., 2024) starcoder2-15b SC-15B
@@
(lla, 2024) Meta-Llama-3-8B-Instruct ML-8B
(gpt, 2024) gpt-3.5-turbo-0125 GPT-3.5
For RQ1 and RQ2, we perform the evaluation using CodeLlama-13B-Instruct (Roziere et al., 2023) as the default LLM, due to its state-of-the-art performance among models of medium size. For RQ3, to evaluate the generalizability of CatCoder, we select a variety of state-of-the-art LLMs with different sizes, including code-specialized models and general-purpose models. The details of the models are listed in Table 2, in which the model IDs are taken from the official releases at HuggingFace(hug, 2024) or the API documentations.
@@ 4.4.3. Environment and Model Access
The experiments are performed on a Linux server with 2 Intel®  Xeon®  Platinum 8358P CPUs and 8 NVIDIA A800 GPUs. For open-source LLMs, we deploy a local API server based on vLLM (Kwon et al., 2023), a unified inference and serving engine for LLMs. The weights of open-source models are available at HuggingFace (hug, 2024). To facilitate future research, we list URLs of the models in our replication package. For closed-source GPT models, we invoke the OpenAI API.
@@ 4.4.4. Configurations
In evaluations, we set $n$ in pass@$k$ to 10 and $k$ to 1, 3, and 5. To control the randomness of the generated code (Chen et al., 2021), we set the temperature to 0.6 and the nucleus sampling (Holtzman et al., 2019) parameter (top_p) to 0.7. In addition, we set the maximum number of generated tokens to 512.
@@ 5.1.1. Results
pass@k (%) pass@k (%)
Method $k=1$ $k=3$ $k=5$ $k=1$ $k=3$ $k=5$
Vanilla 14.92 22.07 25.21 10.78 15.81 18.15
In-File 35.28 43.77 47.58 41.56 55.70 61.60
RepoCoder 40.95 46.91 49.04 49.44 59.47 63.04
CatCoder 44.67 53.97 57.55 52.67 62.00 65.89
The overall result is presented in Table 3. According to Table 3, we observe that the Vanilla baseline performs poorly as expected because the lack of repository context is likely to cause hallucinations. Next, we can conclude that the In-File baseline performs worse than RepoCoder and CatCoder, as in-file context is not enough to provide key information such as cross-file type structures and API usage. In addition, CatCoder outperforms RepoCoder. In terms of the Java benchmark, the pass@$k$ scores increase by up to 17.35%. For the Rust benchmark, the pass@$k$ scores increase by up to 7.14%.
@@ 5.1.2. Qualitative Analysis
To show the advantages of CatCoder over RepoCoder, we also perform a qualitative analysis of the results. Figure 5 presents two cases, in which RepoCoder fails to generate the correct function code while CatCoder succeeds.
@@
Case 2 (method _findWellKnownSimple): The code generated by RepoCoder causes many compilation errors, as it contains non-existent fields and methods in the current scope. In detail, we find that the non-existent CLASS_XXX fields come from another file and RepoCoder retrieves them as context. Without knowing the type context of JavaType and current class (i.e., TypeFactory), the LLM accepts misleading references and generates a non-existent API call stringType().
Answer to RQ1: CatCoder is effective at generating repository-level function code, and it outperforms the RepoCoder baseline by up to 17.35% in terms of pass@$k$ score.

### v2

@@ 3.2.1. Type Context
Figure 3. Dependency Graph and Type context of the triu method in class CMAESOptimizer.
To capture the contextual relationships between types within a code repository, Fig. 3 presents a directed dependency graph $G=(V,E)$, in which each vertex $v\in V$ corresponds to a type defined in the repository (e.g., a class or interface). A directed edge $(A,B)\in E$ is created from type $A$ to type $B$, if $A$ references or depends on $B$ (i.e., $B$ occurs in the body of $A$). The dependency graph is able to identify clusters of contextually related types and provide types that an LLM is likely to use and refer to when generating function code.
In Fig. 3, we build the dependency graph for the triu method in class CMAESOptimizer, starting from its own class. As the graph spans, the vertex set and the edge set will eventually converge, and we can obtain the types that are relevant to CMAESOptimizer. To pass this structural information to the LLM, we create type context from the graph by extracting the fields and function signatures within the types and concatenating them into a string. Specifically, for the string representation, we follow the format proposed by Tufano et al. (tufano2020unit), which has been applied in the field of unit test case generation. The right part of Fig. 3 shows the resulting textual type context for the target method (simplified for illustration).
In large repositories, the full dependency graph can become vast, making the resulting type context too long to fit in an LLM’s context window. However, the relevance between types tends to decrease with the distance between them in the graph. For example, in Fig. 3, programmers working on class CMAESOptimizer are more likely to directly use the API of RealMatrix, a direct neighbor, than the API of RealVector, which is an indirect neighbor at a distance of 2. Therefore, in practice, we constrain the scope of type context creation to the direct neighbors of the initial types in the graph. We also prune types from standard libraries, as we assume LLMs possess general knowledge of these libraries from their pre-training.
@@ 3.2.2. Building Type Context via Static Analyzer
To get detailed information (i.e., fields and methods) of a type from its name, some straightforward heuristics might be used, e.g., locating imports or looking for similar file names. However, these heuristics are not sound and do not work with all statically typed programming languages. Thus, we need to perform an analysis to accurately navigate to a type and find its definitions and implementations. Many language servers serve as static analyzers and provide this type-based analysis, such as clangd (clangd) for C++, Eclipse JDT (eclipsejdtls) for Java, and rust-analyzer (rust_analyzer) for Rust.
Specifically, to build the type context for a function code generation task, we first determine the set of relevant types (as Section 3.2.1 describes). We start with the signature and form a set of types, and the initial element is the type to which the current function belongs. Then we implement a parser that works collaboratively with the static analyzer. The static analyzer provides detailed type information to the parser. The parser then searches for a new occurring type, adds it to the set, and queries the analyzer with its location (i.e., byte offset or line and column offsets in a file). This interactive process continues until all direct neighbors of the initial set of types have been discovered. Finally, as described in Section 3.2.1, we concatenate the fields and method signatures within the set of types and obtain the type context.
@@ 3.3. Code Generation via LLM
Finally, we describe how to utilize the previous results and invoke the LLM to generate code. Fig. 4 illustrates our prompt design, in which there are two fields: CONTEXT and CODE. To construct a prompt, we follow the context format in previous work (zhang2023repocoder; shrivastava2023repository; liu2023repobench), and fill the CONTEXT field by concatenating the type context and the retrieved code chunks. Then we place the docstrings and the signature of the method at the beginning of the CODE field. The prompt serves as input to a frozen LLM [^8] We discuss the details of selected models and model access in Section 4.4., and we expect the model to complete the CODE field by generating the method body (i.e., the green part in Fig. 4). In addition, to post-process the raw output of the LLM, we employ several procedures [^9] The post-processing procedures may vary with different LLMs and programming languages., including truncating texts, prepending function signatures, and removing Markdown syntax.
@@ 4. Experimental Setup
In this section, we first formulate four research questions (RQs). Based on the RQs, we describe the details of the baselines, evaluation metrics, benchmarks, and implementation of our approach.
We aim to answer the following research questions:
  1. RQ1
    How effective is our approach at generating function code in a repository, compared with the baselines? And does our approach generalize to different LLMs?
  2. RQ2
    How do the two components in our approach (i.e., relevant code retrieval and type context extraction) contribute to the overall effectiveness, respectively?
  3. RQ3
@@
  4. RQ4
    Is our approach efficient when generating a large number of functions in large repositories?
In brief, we evaluate the effectiveness of our approach in RQ1, and then conduct an ablation study in RQ2. In RQ3, we study the generalizability of CatCoder and the performance of different models. Finally, in RQ4, we evaluate the scalability of our approach. For implementation and evaluation of effectiveness, we select two statically typed programming languages, Java and Rust, since Java is one of the most widely used programming languages (tiobe), and Rust is a new system-level language that attracts the attention of many developers (rustlang).
@@ 4.1. Baselines
As described in Section 3.3, Fig. 4 shows the prompt template used in our evaluation. As the CONTEXT field is the only variable, we compare our approach with methods that fill different contexts in the prompt. For the repository-level code generation task, there are three types of context: (1) No context, (2) Local context, and (3) Cross-file context. Therefore, we design the following baselines for comparison in RQ1:
@@
    Vanilla: In this method, we prompt LLMs without any repository context. In this case, all natural language descriptions are given, but the CONTEXT field in Fig. 4 is left empty.
@@
    In-File: In this method, we consider code generation as a left-to-right file completion task. We extract code before the function to be generated and put it in the CONTEXT field.
@@
    RepoCoder: RepoCoder (zhang2023repocoder) is an approach that augments the prompt by iteratively retrieving referential code fragments, and it is originally proposed for Python code repositories. We adopt their idea and implement this baseline for Java and Rust based on our retrieval pipeline. Specifically, the initial iteration uses natural language descriptions (i.e., docstrings) as the query for code retrieval. From the second iteration, it uses the previously generated code for retrieval.
4.2. Evaluation Metrics
For the function code generation task, we argue that metrics based on text similarity are not accurate, and it is essential to use metrics based on compilation and execution results. Therefore, we use compilation rate and test passing rate, which evaluate syntax correctness and functional correctness, respectively. Specifically, to calculate such a rate, we follow Chen et al. (humaneval) and use the unbiased score@$k$. It is defined as
score@$k$ = E_Tasks [ 1 - (n-ck)(nk) ], where $n$ is the number of code samples generated per task and $k$ is the parameter in “score@kk” ($n\geq k$). Based on this general formula, we define two detailed metrics that correspond to compilation rate and test passing rate:
  • compile@k. In the formula, $c$ refers to the number of syntactically correct code samples. A code sample is syntactically correct if and only if it passes compiler checks and compiles.
  • pass@k. In the formula, $c$ refers to the number of functionally correct code samples. A code sample is functionally correct if and only if it passes the tests in the benchmark.
We calculate score@$k$ values for each task in the benchmark, and the final score@$k$ for a benchmark is the mean value of all score@$k$ values.
@@ 4.3.1. Java Benchmark
There are several publicly available repository-level benchmarks for Java (ding2024crosscodeeval; yu2024codereval; agrawal2024monitor; zeng2024coderujb). However, it is challenging to use many of them to evaluate CatCoder because they either lack the raw repository code or are not designed for function code generation tasks. Considering the challenges, the CoderUJB (zeng2024coderujb) benchmark is a suitable base benchmark that contains 238 code generation tasks, and we successfully adapt it for our evaluation. In the adaptation process, we filter out a few functions that satisfy one of the following patterns:
@@
    The length of function code exceeds the max output limit (i.e., 512 tokens in our settings).
@@
    The function does not have any modifiers or annotations. This helps us to perform post-processing more easily.
@@
    The function is standalone, which means that it can be moved to a new class [^10] We assume that all Java standard libraries are imported in advance when applying this rule. and still compiles. The removal of standalone functions makes the benchmark more complex and difficult, and helps better evaluate the performance of repository-level code generation.
We apply the filtering rules sequentially: 17 functions are filtered out by the first rule, 5 by the second, and 17 by the third. As a result, after removing these 39 functions, we obtain 199 coding tasks. As the methods in the benchmark are mined by the CoderUJB authors from the Defects4J (just2014defects4j) dataset, it is convenient to evaluate the correctness of the generated code using the existing extensive unit test methods. Specifically, we “patch” the repository with generated code and run the defects4j test -t ... command.
@@ 4.3.2. Rust Benchmark
To our knowledge, the only code generation benchmark for the Rust programming language is MultiPL-E (cassano2023multipl), which translates the original HumanEval (humaneval) and MBPP (mbpp) benchmarks. However, this benchmark is designed for standalone function code generation tasks and thus is not suitable for our evaluation. In this case, we construct a new benchmark named RustEval.
@@
As shown in Table 2, we collect 13 repositories of various categories from crates.io (cratesio), the official Rust package registry. In these repositories, we obtain 90 functions with docstrings that satisfy the official guidelines (rustdoc) (i.e., have detailed descriptions and at least one documentation test). To test the correctness of the generated code, we utilize the documentation tests in the docstrings by running the cargo test --doc ... command.
4.3.3. Benchmark Statistics
Table 3. Statistics of our benchmarks, including number of tasks, average number of lines of code (NLOC) per task, and average cyclomatic complexity number (CCN) per task.
Java Rust
# Tasks 199 90
Avg. NLOC 16.59 10.68
Avg. CCN 4.59 2.64
Table 3 presents the statistics of our benchmarks, where we assess their complexity in terms of the average number of lines of code (NLOC) and cyclomatic complexity number (CCN) per task. Compared to the code completion task in our baseline paper, which requires generating only a single line of code, our function code generation task involves producing more lines and deeper code structures, as reflected in the NLOC and CCN values shown in Table 3. In addition, compared to the Java benchmark, the Rust benchmark is less syntactically verbose and structurally simpler. We attribute this to idiomatic Rust’s reduced boilerplate and more concise abstractions. However, from a semantic standpoint, this does not imply that code generation is easier, as the underlying logic can still be dense and complex.
@@ 4.4.1. Approach
We implement CatCoder for Java and Rust, and the source code is written in Python and Rust. Specifically in Relevant Code Retrieval, we use MPNet V2 (mpnetv2) as the embedding model. It is an open-source model fine-tuned from the base MPNet (song2020mpnet) model using massive training data, including the popular CodeSearchNet (husain2019codesearchnet). For evaluation on the benchmarks, the ground truth code in the repositories is removed during retrieval to prevent leakage. In Type Context Extraction, we use Eclipse JDT.LS (eclipsejdtls) and rust-analyzer (rust_analyzer) as the static analyzers. For Java, our approach interacts with Eclipse JDT.LS through the Language Server Protocol (lsp_spec). For Rust, we use rust-analyzer as a third-party library and implement a binding between Rust and Python.
@@ 4.4.2. Model Selection
(codellama) CodeLlama-7b-Instruct CL-7B
(codellama) CodeLlama-13b-Instruct CL-13B
(deepseekcoder) deepseek-coder-6.7b-instruct DS-6.7B
(codegemma) codegemma-7b-it CG-7B
(codeqwen) CodeQwen1.5-7B-Chat CQ-7B
@@
(llama3) Meta-Llama-3-8B-Instruct ML-8B
For RQ1 and RQ2, we perform the evaluation using CodeLlama-13B-Instruct (codellama) as the default LLM, due to its state-of-the-art performance among models of medium size. For RQ3, to evaluate the generalizability of CatCoder, we select a variety of state-of-the-art LLMs with different sizes, including code-specialized models and general-purpose models. The details of the models are listed in Table 4, in which the model IDs are taken from the official releases at HuggingFace(huggingface) or the API documentations.
@@ 4.4.3. Environment and Model Access
The experiments are performed on a Linux server with capable NVIDIA GPUs. For open-source LLMs, we deploy a local API server based on vLLM (vllm), a unified inference and serving engine for LLMs. The weights of open-source models are available at HuggingFace (huggingface). To facilitate future research, we list URLs of the models in our replication package.
@@ 4.4.4. Configurations
In evaluations, we set $n$ in compile@$k$ and pass@$k$ to 10 and $k$ to 1, 3, and 5. To control the randomness of the generated code (humaneval), we set the temperature to 0.6 and the nucleus sampling (nucleus) parameter (top_p) to 0.7. In addition, we set the maximum number of generated tokens to 512.
@@ 5.1.1. Results
compile@k (%) pass@k (%) compile@k (%) pass@k (%)
Method 11 33 55 11 33 55 11 33 55 11 33 55
Vanilla 33.6 49.4 56.2 14.9 22.1 25.2 18.0 28.3 33.1 10.8 15.8 18.2
In-File 59.7 72.7 76.8 35.3 43.8 47.6 52.1 66.5 70.9 41.6 55.7 61.6
RepoCoder 63.0 74.1 78.2 41.0 46.9 49.0 63.7 73.4 76.4 49.4 59.5 63.0
CatCoder 71.9 84.8 88.4 44.7 54.0 57.5 65.1 75.5 78.9 52.7 62.0 65.9
The overall result is presented in Table 5. According to Table 5, we observe that the Vanilla baseline performs poorly as expected because the lack of repository context is likely to cause hallucinations. Next, we can conclude that the In-File baseline performs worse than RepoCoder and CatCoder, as in-file context is not enough to provide key information such as cross-file type structures and API usage. In addition, CatCoder outperforms RepoCoder. In terms of the Java benchmark, the compile@$k$ and pass@$k$ scores increase by up to 14.44% and 17.35%, respectively. For the Rust benchmark, the compile@$k$ and pass@$k$ scores increase by up to 3.27% and 6.68%.
@@ 5.1.2. Qualitative Analysis
To show the advantages of CatCoder over RepoCoder, we also perform a qualitative analysis of the results. Fig. 5 presents two cases, in which RepoCoder fails to generate the correct function code while CatCoder succeeds.
@@
Case 2 (method _findWellKnownSimple): The code generated by RepoCoder causes many compilation errors, as it contains non-existent fields and methods in the current scope.
In detail, we find that the non-existent CLASS_XXX fields come from another file, and RepoCoder retrieves them as context. Without knowing the type context of JavaType and current class (i.e., TypeFactory), the LLM accepts misleading references and generates a non-existent API call stringType().
5.1.3. Comparison of Java and Rust
As shown in Table 5, the magnitude of improvement provided by CatCoder over the RepoCoder baseline differs between the two languages. On the Java benchmark, CatCoder achieves a relative improvement of up to 17.35% in pass@k, whereas on Rust, the maximum relative improvement is 6.68%. This suggests that CatCoder offers a more substantial advantage in the context of the evaluated Java projects. We attribute this difference to the characteristics of the languages and the structural properties of the benchmark projects. As indicated in Table 3, the Java benchmark is, on average, more structurally complex and verbose than the Rust benchmark. We argue that this increased complexity creates scenarios in which the type context provided by CatCoder becomes a more critical factor for successful code generation, as illustrated in Section 5.1.2.
Table 6. Results of CodeBLEU, using CodeLlama-13B-Instruct. C: CodeBLEU score; N: n-gram match score; W: weighted n-gram match score; S: syntax match score; D: dataflow match score.
Java Rust
Method C N W S D C N W S
Vanilla 0.36 0.23 0.30 0.41 0.37 0.25 0.16 0.21 0.39
In-File 0.51 0.40 0.46 0.58 0.55 0.38 0.29 0.33 0.54
RepoCoder 0.57 0.47 0.53 0.64 0.61 0.46 0.37 0.42 0.60
CatCoder 0.56 0.46 0.52 0.63 0.59 0.44 0.35 0.40 0.60
5.1.4. Semantic Similarity vs. Test Pass Rate
In addition to compile@k and pass@k, we conduct an evaluation using CodeBLEU (ren2020codebleu), a metric that lies between purely text similarity-based and execution-based metrics. Beyond traditional n-gram matching, CodeBLEU incorporates syntactic and semantic matching, thus better capturing the semantic similarity between code snippets. It is computed as: CodeBLEU = α⋅BLEU + β⋅BLEU_weight + γ⋅Match_ast + δ⋅Match_df, where $\mathrm{BLEU}$ is the n-gram match score, $\mathrm{BLEU_{weight}}$ is the weighted n-gram match score, $\mathrm{Match_{ast}}$ is the abstract syntax tree (AST) match score, and $\mathrm{Match_{df}}$ is the dataflow match score. The parameters $\alpha,\beta,\gamma,\text{and }\delta$ are weighting coefficients.
We use an open-source implementation [^11] https://github.com/k4black/codebleu of CodeBLEU, and the results are presented in Table 6. For the Java benchmark, the weights are set to $\alpha=\beta=\gamma=\delta=0.25$. For the Rust benchmark, since the tool cannot extract dataflow from Rust code, we omit the dataflow match component and assign the weights to $\alpha=\beta=\gamma=0.33$ and $\delta=0$.
On both Java and Rust benchmarks, CatCoder and RepoCoder outperform the other two baselines, Vanilla and In-File, across CodeBLEU and all four sub-metrics, suggesting that relevant code retrieval helps LLMs generate code more similar to the ground truth. Notably, the scores of CatCoder and RepoCoder are very close, with RepoCoder even slightly outperforming CatCoder in CodeBLEU. However, since CatCoder achieves higher compile@k and pass@k scores, particularly on the Java benchmark, this indicates that RepoCoder sometimes produces code structurally similar to the ground truth but fails due to subtle issues. This underscores the importance of execution-based metrics in evaluating repository-level code generation.
Answer to RQ1: CatCoder is effective at generating repository-level function code, and it outperforms the RepoCoder baseline by up to 14.44% and 17.35% in terms of compile@$k$ and pass@$k$ scores. Besides, in qualitative analysis, we find that the type context is effective in filtering out misleading reference code.